# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v1

@@ Abstract
We study the covariance model belonging to the Gauss hypergeometric ($\mathcal{GH}$) class, a highly flexible and compactly supported correlation model (kernel). This class includes the well-known Generalized Wendland ($\mathcal{GW}$) and Matérn ($\mathcal{MT}$) kernels as special cases. First, we provide necessary and sufficient conditions for the validity of the $\mathcal{GH}$ model. We then demonstrate that this family of models suffers from identifiability issues under both increasing and fixed-domain asymptotics.
To address this problem, we propose a parsimonious version of the model that adheres to an optimality criterion. This approach results in a new class of compactly supported kernels, which can be seen as an improvement over the $\mathcal{GW}$ model. Additionally, we show that two specific compact support reparameterizations allow us to recover the $\mathcal{MT}$ model, highlighting the advantages and disadvantages of each reparameterization.
[…]
@@ 1 Introduction
Gaussian processes, or Gaussian random fields (RFs), are widely used in various disciplines such as spatial statistics (Cressie,, 1993; Stein,, 1999; Banerjee et al.,, 2004), machine learning (Rasmussen and Williams,, 2006), numerical analysis (Wendland,, 2004), and computer experiments (Santner et al.,, 2003), among others. The Gaussian assumption implies that the finite-dimensional distributions of the RF are fully determined by the mean value and the covariance function. In particular, covariance functions (also known as kernels in the machine learning literature) are positive semidefinite functions that play a fundamental role in modeling relationships between data points in Gaussian RFs, kernel methods, and other machine learning algorithms.
For RFs defined in Euclidean spaces, covariance functions that depend solely on the distance between any pair of points within the domain are called isotropic or radially symmetric. This paper operates under this assumption, which greatly simplifies the discussion. However, isotropic covariances serve as the building blocks for more complex models, such as anisotropic, nonstationary (Paciorek and Schervish,, 2006), multivariate (Gneiting et al.,, 2010), or spatiotemporal extensions (Porcu et al.,, 2020), among others.
To date, there exists a rich catalog of parametric isotropic covariance functions (see, for instance, Rasmussen and Williams, (2006) or Chilès and Delfiner, (2012)), among which the Matérn ($\mathcal{MT}$ hereafter) model is by far the most widely used (see Porcu et al., (2024) for a comprehensive review). It has played a central role in spatial statistics for decades and, more recently, in other disciplines such as numerical analysis, approximation theory, and machine learning. A key advantage of the $\mathcal{MT}$ model is its continuous parameterization of the differentiability of the sample paths of the associated Gaussian RF through a positive smoothness parameter. The larger the smoothness parameter, the higher the level of differentiability of the sample paths. .
While the $\mathcal{MT}$ model remains widely used, particularly in the stochastic partial differential equations approach and its generalizations to non-Euclidean domains (Lindgren et al.,, 2011; Bolin et al.,, 2024), enhancements to the $\mathcal{MT}$ model have been proposed in recent years. These covariance models retain the benefits of the Matérn model while offering additional features (Porcu et al.,, 2024).
Among these, the generalized Wendland ($\mathcal{GW}$ hereafter) (Gneiting,, 2002; Bevilacqua et al.,, 2019) has gained popularity, as it allows for a continuous parameterization of the differentiability of the sample paths of the underlying Gaussian RF, similar to the $\mathcal{MT}$ model, while also being compactly supported. This is a desirable feature from a computational perspective, as it produces sparse covariance matrices (Duff et al.,, 1989; Davis,, 2006).
Several connections between the $\mathcal{MT}$ and $\mathcal{GW}$ models have been established. For instance, Bevilacqua et al., (2019) show that, under specific conditions, these models lead to equivalent Gaussian measures. Additionally, Bevilacqua et al., (2022) demonstrate that the $\mathcal{MT}$ model is a special limit case of a compactly supported reparameterization of the $\mathcal{GW}$ model. Estimation and prediction using $\mathcal{MT}$ and $\mathcal{GW}$ models have been extensively studied under both increasing and fixed-domain asymptotics (Mardia and Marshall,, 1984; Zhang,, 2004; Kaufman and Shaby,, 2013; Bevilacqua et al.,, 2019).
The focus of this paper is the study of the Gauss hypergeometric ($\mathcal{GH}$) covariance, a highly flexible kernel recently proposed by Emery and Alegría, (2022). These authors demonstrate that the $\mathcal{GH}$ model encompasses most well-known covariance models as special or asymptotic cases, including both the $\mathcal{MT}$ and $\mathcal{GW}$ models. In addition, they provide sufficient conditions for the validity of the model. In Porcu et al., (2024), the Gauss hypergeometric ($\mathcal{GH}$) kernel appears to be one of the most promising generalizations of the $\mathcal{MT}$ model. However, the estimation of the $\mathcal{GH}$ covariance model, particularly under fixed-domain asymptotics, has not yet been addressed.
This paper presents several contributions related to the $\mathcal{GH}$ covariance model. First, we provide necessary and sufficient conditions for the validity of the $\mathcal{GH}$ model, offering a complete characterization of the kernel. This result ensures that the model is well-defined and applicable to a wide range of scenarios.
Second, we show that this family of models suffers from identifiability issues under both increasing and fixed-domain asymptotics. These findings highlight a critical limitation of the $\mathcal{GH}$ model in practical applications, particularly when attempting to infer parameters from data. To address this problem, we propose a parsimonious version of the model that adheres to an optimality criterion. This criterion is based on the maximization of the isotropic integral range (Lantuéjoul,, 2002), which can be seen as a measure of the flexibility of the correlation model. For instance, if $d=1$, the maximum integral range corresponds to the maximum area under the unidimensional correlation function.
Using this approach, we derive a new class of compactly supported kernels, which we call the hypergeometric ($\mathcal{H}$ hereafter) model that can be seen as an improvement over the $\mathcal{GW}$ model. The model can also be seen as a generalization of the Euclid’s hat model proposed in Gneiting, (1999), which includes the celebrated spherical models widely used in geostatistics (Chilès and Delfiner,, 2012). Additionally, we show that two specific compact support reparameterizations allow us to recover the $\mathcal{MT}$ model as a special limit case, highlighting the advantages and disadvantages of each reparameterization.
Finally, for a fixed smoothness parameter, we establish strong consistency and the asymptotic distribution of the maximum likelihood (ML) estimator of the microergodic parameter associated with the $\mathcal{H}$ class.
[…]
The $\mathcal{H}$ model has been implemented in the GeoModels package (Bevilacqua et al.,, 2025) for the open-source R statistical environment.
The remainder of this paper is organized as follows. Section 2 provides background material on the $\mathcal{MT}$ and $\mathcal{GW}$ covariance models. Section 3 introduces the $\mathcal{GH}$ model and highlights the identifiability issue. Section 4 first provides necessary and sufficient conditions for the $\mathcal{GH}$ model. Then, following the criterion of maximal integral range, the $\mathcal{H}$ model is proposed. In Section 5, some closed-form expressions for the $\mathcal{H}$ model that do not depend on the computation of special functions are presented. In addition, two compactly supported reparameterizations that include the $\mathcal{MT}$ model as a special case are introduced. In Section 6, we provide the asymptotic properties of the ML estimator of the microergodic parameter. In Section 7, we report a simulation study that explores the finite-sample properties of the ML method when estimating the parameters of the proposed (reparameterized) $\mathcal{H}$ model. In Section 8, we present the analysis of a climate data set. Concluding remarks are provided in Section 9. All proofs have been deferred to Appendix A.
@@ 2 Background: the Matérn and Generalized Wendland correlation models
[…]
Schoenberg, (1938) characterized the class $\Phi_{d}$ as scale mixtures of the characteristic functions of random vectors uniformly distributed on the spherical shell of $\mathbb{R}^{d}$, with any finite and non-negative measure, $F$:
@@
Fourier transforms of radial versions of members of $\Phi_{d}$, for a given $d$, have a simple expression, as reported in Yaglom, (1987) or Stein, (1999). For a member $\phi$ of the class $\Phi_{d}$ such that $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$, we define its isotropic spectral density as
@@
    The Matérn model is a two-parameter globally supported correlation function defined as (Matérn,, 1986):
    ${\cal MT}_{\nu,\alpha}(x)=\begin{cases}1,\quad x=0\\ \frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{\alpha}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{\alpha}\right),\qquad x>0,\end{cases}$ (2)
    where ${\cal K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz and Stegun,, 1970, 9.6.22). The associated spectral density, $\widehat{{\cal MT}}_{\alpha,\nu}$, is given by
@@
    Arguments in Stein, (1999) show that ${\cal MT}_{\nu,\alpha}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $\alpha$ are strictly positive.
    The importance of the Matérn model lies in the parameter $\nu$, which controls the differentiability (in the mean square sense) of the associated Gaussian RF. Specifically, for any integer $l=0,1,\ldots$, a Gaussian RF with a Matérn covariance function is $l$-times mean square differentiable, in any direction, if and only if $\nu>l$. When $\nu=l+1/2$, the Matérn covariance function simplifies into the product of an exponential covariance and a polynomial of order $l$, given by:
    ${\cal MT}_{l+1/2,\alpha}(x)=\exp\left(-\frac{x}{\alpha}\right)\sum_{i=0}^{l}\frac{(l+i)!}{2l!}\binom{l}{i}\left(\frac{2x}{\alpha}\right)^{l-i},\quad l=0,1,\ldots$
@@
    The ${\cal GW}$ model is a three-parameters compactly supported correlation function defined as (Gneiting,, 2002; Bevilacqua et al.,, 2019):
@@
    with $\beta>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and ${}_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun,, 1970, 15.1.1). The associated spectral density is given by:
@@
    For a given $\kappa>-\frac{1}{2}$, a necessary and sufficient condition for ${\cal GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ is (Bevilacqua et al.,, 2024)
    $\mu\geq\mu_{\min}(\kappa,d):=\begin{cases}\frac{\sqrt{8\kappa+9}-1}{2}\text{ if $d=1$ and $-\frac{1}{2}<\kappa<0$}\\ \kappa+\frac{d+1}{2}\text{ otherwise.}\end{cases}$ (6)
@@
    When $\kappa=k\in\mathbb{N}$, the Generalized Wendland correlation simplifies into the product of an Askey (truncated power) correlation with a polynomial of degree $k$:
    ${\cal GW}_{k,\mu,\beta}=\left(1-\frac{x}{\beta}\right)^{\mu+k}_{+}P_{k}(x;\mu,\beta),\quad k=0,1,2,\ldots$ (7)
    where $P_{k}(h,\mu,\beta)$ is a polynomial of degree $k$ that can be expressed as:
    $P_{k}(x;\mu,\beta)=\frac{2^{k}k!}{(2k)!}\sum_{j=0}^{k}c_{j,k}(\mu)\left(\frac{x}{\beta}\right)^{k-j}\left(1-\frac{x}{\beta}\right)^{j},$ (8)
    and $c_{j,k}(\mu)$ are suitable constants (Bevilacqua et al.,, 2024).
    Recently, Bevilacqua et al., (2022) considered a compact support reparameterized version of the ${\cal GW}$ model, and proved that
@@
    where $g(\beta,\mu,\kappa)=\beta\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $[0,\infty)$.
    A simpler compact support reparameterization that leads to the ${\cal MT}$ model as a special limit case has been proposed in Emery et al., (2025):
    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,\mu\beta}(x)={\cal MT}_{\kappa+1/2,\beta}(x),\quad\kappa\geq-0.5,$ (10)
    again with uniform convergence on $[0,\infty)$.
    As a consequence, the two reparameterized ${\cal GW}$ models can be viewed as a generalization of the ${\cal MT}$ model with an additional parameter that, for given smoothness and spatial dependence parameters, allows switching from the world of flexible compactly supported covariance functions to the world of flexible globally supported covariance functions.
@@ 3 The Gauss hypergeometric model and the identifiability problem
The $\mathcal{GH}$ class in Emery and Alegría, (2022) has been defined starting from the following spectral density:
@@
Using some results in Cho and Yun, (2018), Emery and Alegría, (2022) show that, for a given $a>0$, the following conditions:
@@
are sufficient for the positivity of the spectral density on $[0,\infty)$. The associated four-parameter isotropic correlation model is given by:
@@
As in the ${\cal MT}$ and ${\cal GW}$ cases, the ${\cal GH}$ model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\delta$. Furthermore, it is not difficult to see that the ${\cal GH}$ model includes the ${\cal GW}$ and the ${\cal MT}$ models as special cases. For instance,
@@
Although the ${\cal GH}$ model is highly flexible, it suffers from identifiability problems. This is apparent under an increasing domain scenario, where all the parameters can be estimated, since $\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a}$ due to the symmetry of the Gauss hypergeometric function: ${}_{2}F_{1}\left(a;b;\cdot;x\right)={}_{2}F_{1}\left(b;a;\cdot;x\right)$. This implies that there is not a one-to-one correspondence between the parametric space and the correlation function.
To study identifiability under fixed-domain asymptotics, the following theorem establishes the equivalence condition for two distinct Gaussian measures with $\mathcal{GH}$ correlation models that share a common smoothness parameter. Hereafter, we denote by $P(\rho_{i})$, $i=0,1$, a zero-mean Gaussian measure with an isotropic covariance function $\rho_{i}$.
@@
Let $P(\sigma_{i}^{2}{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}})$, $i=0,1$, be two zero-mean Gaussian measures. If $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$, for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$, if and only if
@@
where $L(\beta_{i},\gamma_{i})=\frac{2^{2\delta-d}\Gamma(\beta_{i}-d/2)\Gamma(\gamma_{i}-d/2)}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$ with $i=0,1.$
An immediate consequence of Theorem 1 is that the microergodic parameter of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model is given by $\frac{\sigma^{2}L(\beta,\gamma)}{a^{2\delta-d}}$. Since $L(\beta,\gamma)=L(\gamma,\beta)$, it is clear that there is not a unique correspondence between the parametric space and the microergodic parameter, which implies that the microergodic parameter is not identifiable.
Summarizing our results, the $\mathcal{GH}$ correlation model is highly flexible, including the $\mathcal{GW}$ and the $\mathcal{MT}$ models as special cases. However, it is not identifiable under both increasing and fixed-domain scenarios. To fix this issue, in the next section, we will assume that the second shape parameter ($\gamma$) is a function of the first shape parameter ($\beta$).

### v2

@@ Abstract
We study covariance functions in the Gauss hypergeometric ($\mathcal{GH}$) class, a flexible family that encompasses the Generalized Wendland ($\mathcal{GW}$) and Matérn ($\mathcal{MT}$) models. We derive sharp validity conditions, providing a complete characterization of the admissible parameter space, and show that the model exhibits structural identifiability issues under both increasing- and fixed-domain asymptotics.
To resolve this issue, we introduce a parsimonious compactly supported subclass selected via a maximum integral range criterion. The resulting hypergeometric model can be viewed as a structural refinement of the $\mathcal{GW}$ family and admits compact-support reparameterizations that recover the $\mathcal{MT}$ model as a limit case.
We further establish strong consistency and asymptotic normality of the maximum likelihood estimator of the associated microergodic parameter under fixed-domain asymptotics. Simulation experiments and a real-data application to climate data illustrate the finite-sample behavior and practical performance of the proposed model.
@@ 1 Introduction
Gaussian random fields (RFs) are widely used in spatial statistics, machine learning, numerical analysis, and computer experiments (Cressie, 1993; Stein, 1999; Banerjee et al., 2004; Rasmussen and Williams, 2006; Wendland, 2004; Santner et al., 2003). Under Gaussianity, the dependence structure is fully characterized by the covariance function, whose specification is central for modeling, prediction, and uncertainty quantification.
For RFs on Euclidean spaces, isotropic covariance (kernel) functions— depending only on Euclidean distance—provide a tractable and widely used modeling framework. Among parametric isotropic models, the Matérn ($\mathcal{MT}$) class is arguably the most prominent (Porcu et al., 2024), owing to its smoothness parameter, which continuously controls the mean-square differentiability of the sample paths. A compactly supported alternative is the generalized Wendland ($\mathcal{GW}$) family (Gneiting, 2002; Bevilacqua et al., 2019), which also allows for continuous control of smoothness while yielding sparse covariance matrices and substantial computational gains (Duff et al., 1989; Davis, 2006). Connections between $\mathcal{MT}$ and $\mathcal{GW}$ models have been extensively studied (Bevilacqua et al., 2019, 2022).
We focus on the Gauss hypergeometric ($\mathcal{GH}$) covariance model introduced by Emery and Alegría (2022). This highly flexible class includes both $\mathcal{MT}$ and $\mathcal{GW}$ as special or asymptotic cases and has been recognized as a prominent extension of the Matérn family (Porcu et al., 2024). While sufficient validity conditions were derived in Emery and Alegría (2022), a complete characterization of the parameter space and a detailed inferential analysis remain open.
This paper provides several theoretical and methodological contributions. First, we refine the validity conditions of the $\mathcal{GH}$ model. We show that the sufficient conditions of Emery and Alegría (2022) are also necessary in the parameter range where the two sets of conditions agree. Second, we clarify that the $\mathcal{GH}$ family is not identifiable in the classical parametric sense. The parameter-to-correlation mapping is not injective, and this structural non-identifiability must be distinguished from the well-known inferential limitations under fixed-domain asymptotics, where Gaussian measures may remain equivalent on bounded domains. In such settings, only microergodic parameter combinations are consistently estimable.
To resolve the structural non-identifiability, we introduce an optimality-driven parsimonious subclass of $\mathcal{GH}$. For fixed smoothness and compact support, we select the member that maximizes the isotropic integral range, that is, the total mass of the correlation over $\mathbb{R}^{d}$. Since the integral range coincides with the spectral density at zero frequency, this criterion selects the admissible model with maximal low-frequency spectral power, i.e., the strongest large-scale dependence compatible with the imposed constraints. The integral range provides a global and threshold-free measure of dependence (Vanmarcke, 1983; Lantuéjoul, 2002; Hristopulos and Zukovic, 2011); in dimension $d=1$, it reduces to the area under the correlation function. The proposed selection principle is intrinsic to the model class and does not depend on sampling design or loss functions.
The resulting hypergeometric ($\mathcal{H}$) model constitutes the main contribution of this paper. It provides a canonical and parsimonious representative of the $\mathcal{GH}$ class, preserving compact support, continuous smoothness parameterization, and modeling flexibility. Notably, the $\mathcal{H}$ model satisfies simple and explicit necessary and sufficient validity conditions, reducing to a single lower bound on the shape parameter that is independent of the spatial dimension. In addition, it admits a beta-type scale-mixture representation and may be interpreted as a generalization of the Euclid’s hat model (Gneiting, 1999), encompassing the classical spherical covariance functions widely used in geostatistics (Chilès and Delfiner, 2012).
We further show that a compact-support reparameterization of the $\mathcal{H}$ model yields the Matérn model as a special limit case, thereby extending to the hypergeometric framework the Wendland-to-Matérn connection proved in Bevilacqua et al. (2022).
A simulation study investigates finite-sample behavior under increasing- and fixed-domain scenarios, and a real-data analysis of climate data illustrates that the proposed $\mathcal{H}$ model can outperform the $\mathcal{MT}$ model in terms of model fit, prediction accuracy, and computational efficiency for both Gaussian and Tukey-$h$ RFs (Xu and Genton, 2017).
The $\mathcal{H}$ model is implemented in the GeoModels package (Bevilacqua et al., 2026). The remainder of the paper is organized as follows. Section 2 reviews the $\mathcal{MT}$ and $\mathcal{GW}$ models. Section 3 introduces the $\mathcal{GH}$ model and discusses identifiability. Section 4 provides validity conditions and derives the $\mathcal{H}$ subclass. Section 5 presents closed-form expressions and compact-support reparameterizations. Section 6 establishes asymptotic properties of the microergodic estimator. Sections 7 and 8 report simulation and real-data results. All proofs are collected in Appendix A.
@@ 2 Background: the Matérn and Generalized Wendland correlation models
with $\boldsymbol{h}=\boldsymbol{s}^{\prime}-\boldsymbol{s}$ and $\boldsymbol{s},\boldsymbol{s}^{\prime}\in D$, and $\|\cdot\|$ denoting the Euclidean norm. Gaussian RFs with such covariance functions are called weakly stationary and isotropic.
Schoenberg (1938) characterized the class $\Phi_{d}$ as scale mixtures of the characteristic functions of random vectors uniformly distributed on the spherical shell of $\mathbb{R}^{d}$, with any finite and nonnegative measure, $F$:
@@
Fourier transforms of radial versions of members of $\Phi_{d}$, for a given $d$, have a simple expression, as reported in Yaglom (1987) or Stein (1999). For a member $\phi$ of the class $\Phi_{d}$ such that $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$, we define its isotropic spectral density as
@@
    The Matérn model is a two-parameter globally supported correlation function defined as (Matérn, 1986):
    ${\cal MT}_{\nu,\alpha}(x)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{\alpha}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{\alpha}\right),\qquad x>0,$ (2)
    where ${\cal K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz and Stegun, 1970, 9.6.22). The associated spectral density, $\widehat{{\cal MT}}_{\alpha,\nu}$, is given by
@@
    Arguments in Stein (1999) show that ${\cal MT}_{\nu,\alpha}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $\alpha$ are strictly positive.
    A key feature of the Matérn model is the smoothness parameter $\nu$, which determines the mean-square differentiability of the sample paths of the associated Gaussian RF. In particular, for any integer $l=0,1,\ldots$, the sample paths are $l$-times mean-square differentiable (in any direction) if and only if $\nu>l$. In the special case $\nu=l+\tfrac{1}{2}$, the correlation function reduces to an exponential term multiplied by a polynomial of degree $l$.
@@
    The $\cal GW$} model is a three-parameters compactly supported correlation function defined as (Gneiting, 2002; Bevilacqua et al., 2019):
@@
    with $\beta>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and $_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun, 1970, 15.1.1). The associated spectral density is given by:
@@
    For a given $\kappa>-\frac{1}{2}$, the validity conditions for $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ follow as a special case of Theorem 2 by setting $l=\frac{1}{2}$:
    • (A)
      $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ if and only if $\mu\geq\kappa+\frac{d+1}{2}$, whenever $\kappa\geq 0$ or $d\geq 2$;
    • (B)
      $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{1}$ if $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$, and only if $\mu\geq\kappa+1$, for $\frac{1}{2}<\kappa<0$.
    Since $\kappa+1<\frac{\sqrt{8\kappa+9}-1}{2}$ for all $\frac{1}{2}<\kappa<0$, a gap exists between the necessary and sufficient conditions. This partially corrects Bevilacqua et al. (2024), which claimed in case (B) that $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$ is both necessary and sufficient for $d=1$. Whether this lower bound is also necessary remains an open problem.
@@
    When $\kappa=k\in\mathbb{N}$, the Generalized Wendland correlation simplifies into the product of an Askey (truncated power) correlation with a polynomial of degree $k$.
    Recently, Bevilacqua et al. (2022) considered a compact support reparameterized version of the $\cal GW$} model, and proved that
@@
    where $g(\beta,\mu,\kappa)=\beta\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $0,\infty)$.
    As a consequence, the reparameterized $\cal GW$} model can be viewed as a generalization of the $\cal MT$} model with an additional parameter that, for given smoothness and spatial dependence parameters, allows switching from the world of flexible compactly supported covariance functions to the world of flexible globally supported covariance functions.
@@ 3 The Gauss hypergeometric model and the identifiability problem
The $\mathcal{GH}$ class in Emery and Alegría (2022) has been defined starting from the following spectral density:
@@
Using some results in Cho and Yun (2018), Emery and Alegría (2022) show that, for a given $a>0$, the following conditions:
@@
are sufficient for the positivity of the spectral density on $0,\infty)$. The associated four-parameter isotropic correlation model is given by:
@@
As in the $\cal MT$} and $\cal GW$} cases, the $\cal G$H} model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\delta$.
Furthermore, it is not difficult to see that the $\cal G$H} model includes the $\cal GW$} and the $\cal MT$} models as special cases. For instance,
@@
Although the $\cal G$H} family is highly flexible, it is not identifiable in the usual parametric sense. Indeed, even under increasing-domain asymptotics—where one would in principle expect all parameters to be estimable—the parameter-to-correlation map is not injective. In particular,
$\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a},$
because the Gauss hypergeometric function is symmetric in its first two arguments:
$_{2}F_{1}(a,b;\cdot\,;x)={}_{2}F_{1}(b,a;\cdot\,;x).$
Hence distinct parameter values (obtained by swapping $\beta$ and $\gamma$) induce exactly the same correlation function. This implies that additional constraints are required.
A related inferential limitation arises under fixed-domain (infill) asymptotics, where distinct parameter values may induce equivalent Gaussian measures. The following theorem establishes the equivalence condition for two distinct Gaussian measures with $\mathcal{GH}$ correlation models that share a common smoothness parameter. Hereafter, we denote by $P(\rho_{i})$, $i=0,1$, a zero-mean Gaussian measure with an isotropic covariance function $\rho_{i}$.
@@
Let $P(\sigma_{i}^{2}{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}})$, $i=0,1$, be two zero-mean Gaussian measures and let $\delta>\frac{d}{2}$. If $\frac{d+1}{2}+3\delta<\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0}),$ for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D$, if and only if
@@
where $L(\beta_{i},\gamma_{i})=\dfrac{2^{2\delta-d}\Gamma(\beta_{i}-d/2)\Gamma(\gamma_{i}-d/2)}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$, $i=0,1$.
The restriction to $d=1,2,3$ is typical when studying equivalence of Gaussian measures for isotropic models on $\mathbb{R}^{d}$, and is not specific to the $\cal G$H} class. Indeed, the same dimensional restriction appears in the equivalence results for the $\cal MT$} model (Zhang, 2004), and stems from the integrability conditions required by the criteria of Stein (2004).
An immediate consequence of Theorem 1 is that, for fixed $\delta$, the microergodic parameter is given by
$\frac{\sigma^{2}L(\beta,\gamma)}{a^{2\delta-d}}.$
Because $L(\beta,\gamma)=L(\gamma,\beta)$, this quantity does not distinguish $(\beta,\gamma)$ from $(\gamma,\beta)$. Thus, while the microergodic parameter is the relevant identifiable feature under fixed-domain asymptotics, the individual shape parameters $\beta$ and $\gamma$ are not identifiable without an additional constraint.
Summarizing, the $\mathcal{GH}$ correlation model is highly flexible and includes the $\mathcal{GW}$ and $\mathcal{MT}$ models as special cases. However, it is not identifiable in the usual parametric sense because swapping $(\beta,\gamma)$ leaves the correlation unchanged. This issue can be observed under both increasing and fixed domain scenario. To this end, in the next Section we assume that the second shape parameter ($\gamma$) is a function of the first ($\beta$).